OpenStudy (anonymous):

The ordered pairs (1,36), (2,49), (3,64), and (5,100) represent a function. What is a rule that represents this function? A: y=x^2 B: y=36x C: y=(x+5)^2 D: y=(x+6)^2

OpenStudy (jango_in_dtown):

see that 36=6^2,49=7^2,64=8^2,100=10^2

OpenStudy (jango_in_dtown):

for x=1, f(x)=36=6^2=(1+5)^2 for x=2,f(x)=49=7^2=(2+5)^2 for x=3,f(x)=64=8^2=(3+5)^2 for x=5,f(x)=100=10^2=(5+5)^2

OpenStudy (jango_in_dtown):

so the general form is f(x)=(x+5)^2